40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 9475     Entry time: Sun Dec 15 03:13:15 2013
Author: Den 
Type: Update 
Category: LSC 
Subject: attempt to reduce carm offset 

X,Y arms were stabilized using ALS and moved 5 nm from the resonance, PRMI was locked on sideband using REFL165 I&Q. POP angular servo was running; PRMI RIN was good (~2-3%)

During slow offset reduction I was sweeping MICH, PRCL and POP servos for instabilities due to possible optical gain variations, loops were fine.

I could reduce offset down to ~200 pm and then lost lock due to 60 Hz oscillations as shown on the attached plot "arm_offset"

Arms were stabilized with RMS comparable to the offset and power in arms was fluctuating from 3 to 45.

60 Hz line most probably comes from MICH. RMS is dominated by the power lines and is ~ 1 nm as seen on the plot "PRMI_CAL". I think this is too much but we need to do simulations.

Attachment 1: ARM_OFFSET.pdf  93 kB  | Hide | Hide all
ARM_OFFSET.pdf
Attachment 2: PRMI_CAL.pdf  68 kB  | Hide | Hide all
PRMI_CAL.pdf
ELOG V3.1.3-